THE S_{1} - S_{0} FLUORESCENCE EXCITATION SPECTRUM OF FORMIC ACID

\underline{L.M. BEATY}, D.C. MOULE, Department of Chemistry, Brock University, St. Catharines, Ontario L2S 3A1; R.H. JUDGE, Department of Chemistry, University of Wisconsin-Parkside, Kenoska, WI 5341-2000; H. LIU and E.C. LIM, Knight Chemical Laboratory, University of Akron, Akron, OH 44325-3601.

The fluorescence excitation spectrum of formic acid monomer (HCOOH) has been recorded under jet-cooled conditions in the region, 268-257 nm. A rotational analysis of the first four bands was used to establish the allowed and forbidden vibronic character of the transition as well as the torsional level splittings. A simulation of the torsional progression was carried out by ab-initio calculations at differing levels of approximation as a starting point for the refinement of the torsional and wagging potential functions.